2017 2016 2015 2014 2013
Sales $ 688,869 $ 459,246 $ 361,611 $ 245,160 $ 181,600
Cost of goods sold 353,773 235,915 187,622 126,682 92,616
Accounts receivable 33,341 26,820 24,843 14,317 12,440


Compute trend percents for the above accounts, using 2013 as the base year.
Answer #1

Trend percent for net sales

Choose numerator / choose denominator =
Analysis period net sales / base year net sales = sales
2017 688869 / 181600 = 379%
2016 459246 / 181600 = 253%
2015 361611 / 181600 = 199%
2014 245160 / 181600 = 135%
2013 181600 / 181600 = 100%

Trend percent for cost of goods sold

Choose numerator / choose Denominator
Analysis period cost of goods sold / base year cost of goods sold = cost of goods sold
2017 353773 / 92616 = 382%
2016 235915 / 92616 = 255%
2015 187622 / 92616 = 203%
2014 126682 / 92616 = 137%
2013 92616 / 92616 = 100%

Trend percent for accounts receivable

Choose numerator / choose denominator =
Analysis period accounts receivable / base year accounts receivable = accounts receivable
2017 33341 / 12440 = 268%
2016 26820 / 12440 = 216%
2015 24843 / 12440 = 200%
2014 14317 / 12440 = 115%
2013 12440 / 12440 = 100%